If the Cesium Co nucleus had 78 55 Protons and neutrons; then its ? radius would be
Answer #1

total  nucleon of Cs = protons + neutrons

total  nucleon of Cs (A) = 55 + 78 = 133

volume of single nucleon = \frac{4}{3}\pi R_{0}^{3}

where Ro is radius of single nucleon . it is around in the range of 1\times 10^{-15} m

volume of Cs nucleon =\frac{4}{3}\pi R^{3}

where R is radius of Cs nucleon

so the volume of the Cs nucleus = number of Cs nucleon \times   volume of single nucleon

\frac{4}{3}\pi R^{3} = A\times \frac{4}{3}\pi R_{0}^{3}

R^{3} =A\times R_{0}^{3}

R =A^{\frac{1}{3}}\times R_{0}

R =133^{\frac{1}{3}}\times 1\times 10^{-15}

R =5.1\times 10^{-15}m